OpenStudy (anonymous):

PLEASE HELP!! Solve for x: (3/x)-(2/3)=5/6

OpenStudy (anonymous):

move -2/3 onto left side it will become+2/3 then take lcm

OpenStudy (moongazer):

apply the properties of equality Ax+B=C Ax+B+(-B)=C-B Ax=C-B (1/A)(Ax)=(C-B)(1/A) X=(C-B)/A

OpenStudy (anonymous):

I dont undertstand! :/

OpenStudy (anonymous):

3/x=5/6+2/3 lets try next

OpenStudy (anonymous):

ok so i worked through the problem to where its 9/6=3/x and so i cross multiplyed and got x=2 ... is that right!?

OpenStudy (anonymous):

@sail1234

OpenStudy (anonymous):

lcm of 5/6+2/3 = 5+2*(2)/6 =9/6 equalise 3/x=9/6 reverse on both sides x/3= 6/9 can u proceed

OpenStudy (anonymous):

then you cross multiply?? it would still be 2... :/

OpenStudy (anonymous):

@sail1234

OpenStudy (moongazer):

x=2 is correct :)

OpenStudy (anonymous):

ok thanks!! ;)

OpenStudy (anonymous):

yes well done

OpenStudy (moongazer):

@cassie4math try checking www.khanacademy.org I think it has many videos about linear equations :)